OpenStudy (anonymous):

a triangle with sides of lengths 16,30,and 35 is a right triangle? a) true b)false

OpenStudy (jadeishere):

can it fit into this situation? 16^2 + 30^2 = 35^2 if it does, then yes, it's a right triangle, but if not, it is not a right triangle

OpenStudy (anonymous):

False. Use the Pythagorean Theorem: In a right triangle, a² + b² = c² 16² + 30² = 35² 256 + 900 = 1225 1156 ≠ 1225 Thus the triangle is not a right triangle

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome :)